OpenStudy (alissad):

Which of the following expressions is equivalent to the expression below? log 7x4

OpenStudy (legomyego180):

http://www.purplemath.com/modules/logrules.htm

OpenStudy (alissad):

A. log 7 - 4 log x B. 4 log 7 + log x C. 4 log 7 - log x D. log 7 + 4 log x

OpenStudy (alissad):

those are the answers it gives

OpenStudy (legomyego180):

I think its B but I would double check. Its been a while since Ive done this.

OpenStudy (legomyego180):

this is \[\log 7x^4\] right?

OpenStudy (alissad):

yeah

OpenStudy (legomyego180):

The correct answer is D

OpenStudy (legomyego180):

Review and memorize log rules. \[ \log x^4 = 4 \log x\] You can rewrite exponets in front

OpenStudy (legomyego180):

But you have to keep whatever was raised to the power attached, which is important in this problem. What I mean is that it has to be 4 log x because the x was raised to the 4. It cant be 4 log 7 by itself.